Factorise the following expressions.
p^2 +6p–16 ????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

p^2+6p–16
We observed that, -16 = -2×8 and 8+(-2) = 6
p^2 + 6p –16
= p^2–2p+8p–16
= p(p–2)+8(p–2)
= (p+8)(p–2)
So, p^2+6p–16 = (p+8)(p–2)

Thanks